# Data-Driven Computation of Polytopic Invariant Sets for Noisy Nonlinear Systems

### v1

@@ 4. Data-Driven Computation of Robust Polytopic Control Invariant Set
A finite number of bounded-degree inequality and equality constraints defines a polytope [putinar1993positive]. The polytopic set of all plants and noise values consistent with the data dictionary $\mathcal{D}=\left\{x_{k},u_{k},x_{k+1}\}_{k=0}^{T-1}\right$. and known noise upper bound $\epsilon$ is the consistency set which is defined as follows
(4) $\mathcal{P}_{1}\doteq\{\Theta\in\mathbb{R}^{n\times d_{f}}:\|x_{k+1}-\Theta\phi(x_{k},u_{k})\|_{\infty}\leq\epsilon\}.$
@@
(5) $\mathcal{P}_{1}=\left\{\theta:\begin{bmatrix}A\\ -A\end{bmatrix}\theta\leq\begin{bmatrix}\epsilon\mathbf{1}+\xi\\ \epsilon\mathbf{1}-\xi\end{bmatrix}\right\}$.
@@
(6) $A\doteq\begin{bmatrix}I\otimes\phi^{T}(x_{0},u_{0})\\ \vdots\\ I\otimes\phi^{T}(x_{T-1},u_{T-1})\end{bmatrix},\quad\xi\doteq\begin{bmatrix}x_{1}\\ \vdots\\ x_{T}\end{bmatrix}$.
The main reason for expressing the consistency set in this polytopic form is computational. This representation allows the set of all possible system parameters to be finitely characterized by its vertices.
@@
By Assumption 4, the consistency is defined by the convex hull of its vertices. These vertices can be computed using standard vertex enumeration algorithms[avis1991pivoting]. We denote this set of all computed vertices of consistency set to be $Q=\{\theta^{*,1},\theta^{*,2},\dots,\theta^{*,q}\}$ that contains $q$ vertices in total.
The dependency of the function $F$ represented in (2) on the system parameters is made explicit through the following sequence of relationships. First, each vertex of the consistency set represents a distinct possible model for the unknown noisy, nonlinear systems. Letting $\Theta^{*,i}$ be the matrix form of the vectorized vertex $\theta^{*,i}\forall i\in N_{q}$, this model is given by:
(7) $f(x_{k},u_{k};\theta^{*,i})=\Theta^{*,i}\phi(x_{k},u_{k}),\forall i\in N_{q},$
where $f(x_{k},u_{k};\theta^{*,i})$ denotes the system’s dynamic associated with the vectorized parameter $\theta^{*,i}$. As Assumption 2 holds, the convex components of $f$, $g$ and $h$ inherit this dependency, such that for each vertex we have the decomposition:
(8) $f(x_{k},u_{k};\theta^{*,i})=g_{i}(x_{k},u_{k})-h_{i}(x_{k},u_{k}),\forall i\in N_{q},$
where $g_{i},h_{i}$ represents vector of convex function associated with $i^{th}$ component of the set $Q$.
@@
Since the basis functions $\phi$ that constitute $f$ are known, this decomposition into convex components $g$ and $h$ can be performed using established methods, such as those in [niu2018difference]. It is important to note, however, that this DC decomposition is not unique. This non-uniqueness has a direct impact on the results, as different choices for $g,h$ can lead to different levels of conservatism in the final calculation of the $\lambda$-contractive set. The reason for this is that the structure of the function $F$ is dependent on the chosen decomposition, which in turn alters the feasible set of the final optimization problem to be solved.
Consequently, the function $F$, which is constructed from these DC components and their linearizations, is also parameterized by the specific vertex $\theta^{*,i}\in Q$ under consideration. This leads to the following formal expression:
@@
Let $\theta^{*,i}$ be the $i^{th}$ vertex of the consistency set, each represents a possible parameter of the nonlinear system. For any vector $c\in\mathbb{R}^{n}$, we define the function $F(x,u,c;\theta^{*,i})$ as:
(9) $F(x,u,c;\theta^{*,i})=$ $\sum_{j\in b_{+}}c_{j}\left(g_{j}(x,u;\theta^{*,i})-h_{j}^{L}(x,u;\theta^{*,i})\right)$
$+\sum_{j\in b_{-}}c_{j}\left(g_{j}^{L}(x,u;\theta^{*,i})-h_{j}(x,u;\theta^{*,i})\right),$
where $g_{j}(\cdot,\cdot;\theta^{*,i})$ and $h_{j}(\cdot,\cdot;\theta^{*,i})$ are the $j^{th}$ convex components of the system dynamics corresponding to the vertex $\theta^{*,i}$, and $g_{j}^{L}$, $h_{j}^{L}$ are their respective linearizations at the origin. The index sets defined as $b_{+}=b_{+}(c)=\{j\in\mathbb{N}_{n}:c_{j}\geq 0\}$ and $b_{-}=b_{-}(c)=\{j\in\mathbb{N}_{n}:c_{j}<0\}$
Having redefined the function $F$ to account for the parameter uncertainty, we now propose the following sufficient condition for a set to be $\lambda$-contractive for the unknown noisy nonlinear system.
@@
Let Assumptions 14 hold. Consider a polytope $\Omega=\{x\in\mathbb{R}^{n}:Hx\leq\mathbf{1}\}\subseteq X$, and let $Q=\{\theta^{*,1},\dots,\theta^{*,q}\}$ be the set of vertices of the consistency set. If there exist control actions $u^{p}=u^{p}_{v}\in U$ defined for each vertex $v^{p}$ of $\Omega$ (for $p\in\mathbb{N}_{n_{v}}$), such that
(10) $F(v^{p},u^{p}_{v},H_{i}^{T};\theta^{*,i})\leq\lambda_{w}-\phi_{W}(H_{i}^{T}),\forall\theta^{*,i}\in Q$
@@
holds for all polytope vertices $p\in\mathbb{N}_{n_{v}}$, all polytope facets $i\in\mathbb{N}_{n_{h}}$, then for a given $\lambda_{w}\in[0,1]$, the polytope $\Omega$ is a $\lambda$-contractive set for all system compatible with data dictionary $\mathcal{D}$.
Proof.   Under Assumption 4, the consistency set is a compact polytope. This implies for any $\theta$ within this set, we can express $\theta$ as a convex combination of the set’s vertices, such that $\theta=\sum_{j=1}^{q}\beta_{j}\theta^{*,j}$, with coefficients $\beta_{j}\geq 0$ and $\sum_{j=1}^{q}\beta_{j}=1$. Therefore, each vertex of the consistency set is interpreted as a ”worst-case” realization of the unknown system parameters. Now, we can write the dynamic associated with vector parameter $\theta$ within the consistency set as follows:
(11) $f(x_{k},u_{k};\theta)=$ $\sum_{j=1}^{q}\beta_{j}f(x_{k},u_{k};\theta^{*,j})$
$=$ $\sum_{j=1}^{q}\beta_{j}(g_{j}(x_{k},u_{k};\theta^{*,j})-h_{j}(x_{k},u_{k};\theta^{*,j}))$
$=$ $\sum_{j=1}^{q}\beta_{j}g_{j}(x_{k},u_{k};\theta^{*,j})-\sum_{j=1}^{q}\beta_{j}h_{j}(x_{k},u_{k};\theta^{*,j})$
where $g_{j},h_{j}$ are convex components of the systems dynamics associated with the $j^{th}$ vertex. Let (11) be the decomposition for system corresponding to $j^{th}$ vertex in consistency set and the definition of $F(x,u,c;\theta^{*,i})$ proposed in (9), we can write:
(12) $F(x,u,c;\theta)=$ $\sum_{j=1}^{q}\beta_{j}F(x,u,c;\theta^{*,j})$
Based on Property (1), the function $F$ is convex. Considering the initial polytope $\Omega=\{x\in\mathbb{R}^{n}:Hx\leq\mathbf{1}\}\subseteq X$, we arrive at the following relationship:
(13) $F(v^{p},u^{p}_{v},H_{i}^{T};\theta)=\sum_{j=1}^{q}\beta_{j}F(v^{p},u^{p}_{v},H_{i}^{T};\theta^{*,j}).$
where $v^{p}$ denotes the vertex of $\Omega$, and $u^{p}_{v}$ represents the input of the system at $v^{p}$ vertex. We can now adapt the sufficient condition for $\lambda$-contractiveness, originally proposed for known nonlinear systems in [fiacchini2010computation] (Property 3). It can be said that each vertex $\theta^{*,j}$ is associated with a worst case system parameters in consistency set and if for all $\beta_{j}$ in (12), the condition (3) holds, the set $\Omega$ will be $\lambda$-contractive for all systems compatible with the available information.
It remains to show that there exists a control law $u$ such that for any initial state $x_{0}\in\Omega$, the trajectory of the system (1) subject to process noise satisfies $x_{k}\in\lambda_{w}^{k}\Omega$ for all $k\geq 0$. For the second part of the theorem where we need to prove there exists a control law for any initial state within the polytope, we explicitly construct the control law for any state $x\in\Omega$. Since $\Omega$ is a polytope, any $x$ can be expressed as a convex combination of its vertices $\{v^{1},\dots,v^{n_{v}}\}$:
$x=\sum_{p=1}^{n_{v}}\alpha_{p}v^{p},\quad\text{with }\alpha_{p}\geq 0,\sum_{p=1}^{n_{v}}\alpha_{p}=1.$
Let $u^{p}_{v}$ be the control actions at the vertices that satisfy the condition in Theorem 1. We define the control law $u\in U$ using the same convex combination coefficients $\alpha_{p}$:
$u=\sum_{p=1}^{n_{v}}\alpha_{p}u^{p}_{v}.$
Next, we must show that for any $x_{0}\in\Omega$ and any disturbance $w_{0}\in W$, the successor state $x_{k+1}=f(x_{k},u_{k};\theta)+w_{k}$ is contained in $\lambda_{w}\Omega$ for all $\theta$ in the consistency set $\mathcal{P}_{1}$.
From the properties of the support function and the function $F$, we have the following upper bound for all $\theta$, and $\forall i\in N_{n_{h}}$:
(14) $H_{i}(f(x_{k},u_{k};\theta)+w_{k})\leq F(x_{k},u_{k},H_{i};\theta)+\phi_{W}(H_{i})$
As established, the function $F(x,u,c;\theta)$ is convex in the pair $(x,u)$. Using (13) and as shown in the proof of Theorem 1, the condition $F(v^{p},u^{p}_{v},H_{i}^{T};\theta)\leq\lambda_{w}-\phi_{W}(H_{i}^{T})$ holds for all $\theta$ in consistency set. We can write:
$H_{i}^{T}x_{k+1}$ $\leq\left(\sum_{p=1}^{n_{v}}\alpha_{p}F(v^{p},u^{p}_{v},H_{i}^{T};\theta)\right)+\phi_{W}(H_{i}^{T})$
$\leq\left(\sum_{p=1}^{n_{v}}\alpha_{p}(\lambda_{w}-\phi_{W}(H_{i}^{T}))\right)+\phi_{W}(H_{i}^{T})$
$=(\lambda_{w}-\phi_{W}(H_{i}^{T}))\sum_{p=1}^{n_{v}}\alpha_{p}+\phi_{W}(H_{i}^{T})$
$=\lambda_{w}-\phi_{W}(H_{i}^{T})+\phi_{W}(H_{i}^{T})=\lambda_{w}.$
Since $H_{i}^{T}x_{k+1}\leq\lambda_{w}$ for all facets $i\in N_{n_{h}}$ and vertices of consistency set, we have shown that $x_{k+1}\in\lambda_{w}\Omega$. $\Box$
@@
Based on Theorem 1, we now present Algorithm 1. This optimization-based procedure is designed to compute the largest possible polytopic $\lambda$-contractive set within the state constraints for the unknown noisy nonlinear system.
Algorithm 1 Data-Driven Computation of a $\lambda$-Contractive Set for Noisy Nonlinear Systems
1:Given $\Omega$, $Q=\{\theta^{*,1},\dots,\theta^{*,q}\}$, and $\epsilon$
2:for $s=1,\dots,q$ do
@@
4:   solve for $\alpha^{s,p}$ and $\forall i\in N_{n_{h}}$:
5:    $\alpha^{s,p}=\max_{\gamma^{p}>0,u^{p}\in U}\gamma^{p}$
6:   s.t. $F(\gamma^{p}v^{p},u^{p},H_{i}^{T};\theta^{*,s})\leq\lambda_{w}\gamma^{p}-\phi_{W}(H_{i}^{T})$
@@
10:return $\alpha\Omega$
@@
The main computational challenge of Algorithm 1 lies in the nested loops, which require solving a convex optimization problem for each pair of vertices from the consistency set $Q$ and the candidate polytope $\Omega$. The total number of optimization problems to be solved is $q\times n_{v}$.
4.1. Randomized Method for Enlarging $\lambda$-Contractive Set
The method presented in Algorithm 1 is effective at finding the largest $\lambda$-contractive set for a given initial polytope shape. However, the final result is inherently limited by this initial choice and may be a conservative approximation of the true maximal invariant set. To overcome this limitation and find a larger, less conservative set, an iterative enlarging method can be applied. Such a method systematically expands a known invariant set by adding new, provably safe points within the state constraint set. In this subsection, we review an enlarging procedure based on the work of [fiacchini2010computation].
@@ 4.1. Randomized Method for Enlarging $\lambda$-Contractive Set
The authors in [fiacchini2010computation] show that an existing $\lambda$-contractive set can be enlarged to reduce conservatism. For low-dimensional systems, they suggest that one can generate random candidate points in the state space and test if they can be added to the set as mentioned in Remark 6. However, they note that this approach becomes inefficient and non-trivial for high-dimensional systems. To address this, they propose an alternative, more systematic procedure which involves solving the following convex optimization problem:
(15) $\max_{\hat{x}\in X,\hat{u}\in U}$ $c^{T}\hat{x}$
s.t. $F(v^{p},u^{p}_{v},H_{i}^{T};\theta^{*,j})\leq\lambda_{w}-\phi_{W}(H_{i}^{T}),$
$i\in\mathbb{N}_{n_{h}},\forall p\in\mathbb{N}_{n_{v}},\forall j\in N_{q}$
for all to find a new point guaranteed to expand the known invariant region.
This section has presented a complete framework for computing robust $\lambda$-contractive sets for unknown nonlinear systems from data. The proposed algorithm provides a systematic procedure for this task. In the following section, we will demonstrate the effectiveness of this approach with a numerical example.
5. Simulation Results
To illustrate the effectiveness of the proposed data-driven method, we apply it to a numerical example. We consider the discretized system dynamics presented in [cannon2003nonlinear]. The data dictionary, $\mathcal{D}=\{(x_{k},u_{k},x_{k+1})\}_{k=0}^{k=30}$, are generated by using the system dynamics given by the equation:
(16) $x_{k+1}=\begin{bmatrix}1&\alpha\\ \alpha&1\end{bmatrix}x_{k}+\alpha\left\{\mu\begin{bmatrix}1\\ 1\end{bmatrix}+(1-\mu)\begin{bmatrix}1&0\\ 0&-4\end{bmatrix}x_{k}\right\}u_{k}$

### v2

@@ IV Data-Driven Computation of Robust Contractive Set
The feasible parameter set of all systems compatible with the data dictionary $\mathcal{D}=\left\{x_{k},u_{k},x_{k+1}\}_{k=0}^{T-1}\right$. and the known noise upper bound $\epsilon$ is the consistency set which is defined as follows
$\mathcal{P}_{1}\doteq\Big\{\Theta\in\mathbb{R}^{n\times d_{f}}:$ $\|x_{k+1}-\Theta\phi(x_{k},u_{k})\|_{\infty}\leq\epsilon,$ (4)
$\forall k\in\{0,\dots,T-1\}\Big\}.$
@@
$\mathcal{P}_{1}=\left\{\theta:\begin{bmatrix}A\\ -A\end{bmatrix}\theta\leq\begin{bmatrix}\epsilon\mathbf{1}+\xi\\ \epsilon\mathbf{1}-\xi\end{bmatrix}\right\}$, (5)
@@
$A\doteq\begin{bmatrix}I\otimes\phi^{T}(x_{0},u_{0})\\ \vdots\\ I\otimes\phi^{T}(x_{T-1},u_{T-1})\end{bmatrix},\quad\xi\doteq\begin{bmatrix}x_{1}\\ \vdots\\ x_{T}\end{bmatrix}$, (6)
The main reason for expressing the consistency set in this polytopic form is computational. The polytopic representation of the consistency set for linear systems is investigated in [14]. This representation allows the set of all possible system parameters to be finitely characterized by its vertices.
@@
By Assumption 4, the consistency set is defined by the convex hull of its vertices. These vertices can be computed using standard vertex enumeration algorithms[1]. We denote the set of all computed vertices of the consistency set to be $Q=\{\theta^{*,1},\theta^{*,2},\dots,\theta^{*,n_{q}}\}$ that contains $n_{q}$ vertices in total.
We now explicitly characterize the dependence of the function F in (2) on the system parameters through the following relationships: Each computed vertex of the consistency set represents a distinct possible model parameter for the noisy nonlinear system. For each vertex $s\in\mathbb{N}_{n_{q}}$, we define $\Theta^{*,s}$ as the matrix representation of $\theta^{*,s}$. The associated dynamic is:
$f(x_{k},u_{k};\theta^{*,s})=\Theta^{*,s}\phi(x_{k},u_{k}),\forall s\in\mathbb{N}_{n_{q}},$ (7)
Since Assumption 2 holds, the convex components of $f$, $g$ and $h$, inherit this dependency, such that for each vertex we have the decomposition:
$f(x_{k},u_{k};\theta^{*,s})=g(x_{k},u_{k};\theta^{*,s})-h(x_{k},u_{k};\theta^{*,s}),\forall s\in\mathbb{N}_{n_{q}}$ (8)
where $g(\cdot;\theta^{*,s}),h(\cdot;\theta^{*,s})$ represents vector of convex function associated with $s^{th}$ component of the set $Q$.
@@
Because the basis functions $\phi$ are known, decomposing $f$ into convex components $g$ and $h$ is computationally tractable using established methods like the DC Sums-of-Squares (DC-SOS) framework [12], which leverages spectral decomposition and recursive monomial factorization.
Consequently, the function $F$, which is constructed from these DC components and their linearizations, is also parameterized by the specific vertex $\theta^{*,s}\in Q$ under consideration. This leads to the following formal expression:
@@
For any vector $c\in\mathbb{R}^{n}$, we define the function $F(x,u,c;\theta^{*,s})$ as:
$F(x,u,c;\theta^{*,s})=$ $\sum_{j\in b_{+}}c_{j}\left(g_{j}(x,u;\theta^{*,s})-h_{j}^{L}(x,u;\theta^{*,s})\right)$ (9)
$+\sum_{j\in b_{-}}c_{j}\left(g_{j}^{L}(x,u;\theta^{*,s})-h_{j}(x,u;\theta^{*,s})\right),$
where $g_{j}(\cdot,\cdot;\theta^{*,s})$ and $h_{j}(\cdot,\cdot;\theta^{*,s})$ are the $j^{th}$ convex components of the system dynamics corresponding to the vertex $\theta^{*,s}$ where $s\in\mathbb{N}_{n_{q}}$, and $g_{j}^{L}$, $h_{j}^{L}$ are their respective linearizations at the origin. The index sets are defined as $b_{+}=b_{+}(c)=\{j\in\mathbb{N}_{n}:c_{j}\geq 0\}$ and $b_{-}=b_{-}(c)=\{j\in\mathbb{N}_{n}:c_{j}<0\}$
Having redefined the function $F$ to account for the parameter uncertainty, we now propose the following sufficient condition for a set to be robust $\lambda_{\omega}$-contractive for the unknown noisy nonlinear system.
@@
Let Assumptions 14 hold. Consider a polytope $\Omega=\{x\in\mathbb{R}^{n}:Hx\leq\mathbf{1}\}\subseteq X$. If there exist control actions $u^{p}=u^{p}_{v}\in U$ defined for each vertex $v^{p}$ of $\Omega$, such that
$F(v^{p},u^{p}_{v},H_{i}^{T};\theta^{*,s})\leq\lambda_{\omega}-\phi_{W}(H_{i}^{T}),\forall s\in\mathbb{N}_{n_{q}},$ (10)
@@
holds for all polytope facets $i\in\mathbb{N}_{n_{h}}$, then for a given $\lambda_{\omega}\in[0,1]$, the polytope $\Omega$ is a $\lambda_{\omega}$-contractive set for all systems compatible with data dictionary $\mathcal{D}$.
Proof.
Please check the Appendix. ∎
@@
Based on Theorem 1, we now present Algorithm 1. This optimization-based procedure is designed to compute the largest possible polytopic $\lambda_{\omega}$-contractive set within the state constraints for the unknown noisy nonlinear system.
Algorithm 1 Computation of a $\lambda_{\omega}$-Contractive Set
1:Given $\Omega_{0}$, $Q=\{\theta^{*,1},\dots,\theta^{*,n_{q}}\}$, $\lambda_{\omega}$,
2:for $s=1,\dots,n_{q}$ do
@@
4:   solve the following for $\alpha^{s,p}$ and $\forall i\in\mathbb{N}_{n_{h}}$:
5:    $\alpha^{s,p}=\max\limits_{\gamma^{p}>0,u^{p}\in U}\gamma^{p}$
6:   s.t. $F(\gamma^{p}v^{p},u^{p},H_{i}^{T};\theta^{*,s})\leq\lambda_{\omega}\gamma^{p}-\phi_{W}(H_{i}^{T})$
@@
10:return $\alpha\Omega_{0}$
It is worth mentioning that the variable $\gamma^{p}$ acts as a scalar multiplier determining the maximum permissible scaling of each candidate vertex $v^{p}$ along its original direction.
@@
The computational complexity involves two stages: First, enumerating the $n_{q}$ vertices of the consistency set from $2T$ inequalities scales exponentially with the parameter dimension. Second, the optimization in Algorithm 1 scales proportionally to $n_{q}\times n_{v}\times n_{h}$. Therefore, selecting the initial candidate polytope $\Omega_{0}$ presents a clear trade-off: increasing its complexity yields a less conservative $\lambda$-contractive set, but directly increases the computational load.
Remark 6.
The DC decomposition (f=ghf=g-h) relaxes the intractable non-convex invariance condition into a tractable convex optimization problem by constructing a convex upper bound, $F$. Crucially, this decomposition is not unique. This flexibility can be exploited to reduce conservatism: by evaluating multiple valid decompositions for a fixed initial polytope $\Omega_{0}$, one can select the decomposition that maximizes the scaling factor $\alpha$, resulting in the least conservative $\lambda$-contractive set.
IV-A Enlarging $\lambda$-Contractive Set
The method presented in Algorithm 1 is effective at finding the largest $\lambda$-contractive set for a given initial polytope shape. However, the final result is inherently limited by this initial choice and may be a conservative approximation of the true maximal invariant set. To overcome this limitation and find a larger, less conservative set, an iterative enlarging method can be applied. Such a method systematically expands a known invariant set by adding new, provably safe points within the state constraint set. In this subsection, we review an enlarging procedure based on the work of [8].
@@ IV-A Enlarging $\lambda$-Contractive Set
The authors in [8] show that an existing $\lambda$-contractive set can be enlarged to reduce conservatism. For low-dimensional systems, they suggest that one can generate random candidate points in the state space and test if they can be added to the set as mentioned in Remark 7. However, this approach becomes inefficient and non-trivial for high-dimensional systems. To address this, an alternative, more systematic procedure involves solving the following convex optimization problem to determine a point $\hat{x}\in X$ to enlarge $\Omega=\{x\in\mathbb{R}^{n}:Hx\leq\mathbf{1}\}$. Given $c\in\mathbb{R}^{n}$
$\max_{\hat{x}\in X,\hat{u}\in U}$ $c^{T}\hat{x}$ (11)
s.t. $F(\hat{x},\hat{u},H_{i}^{T};\theta^{*,s})\leq\lambda_{\omega}-\phi_{W}(H_{i}^{T}),$
$i\in\mathbb{N}_{n_{h}},\forall s\in\mathbb{N}_{n_{q}},$
[…]
V Simulation Results
In this section, we illustrate the effectiveness of our proposed method to compute the largest possible invariant set within the state constraints for noisy nonlinear systems in a data-driven manner.
Case Study: We consider the discretized system dynamics presented in [6] given by the equation:
$x_{k+1}$ $=\begin{bmatrix}1&\tau\\ \tau&1\end{bmatrix}x_{k}+\tau\left\{\mu\begin{bmatrix}1\\ 1\end{bmatrix}+(1-\mu)\begin{bmatrix}1&0\\ 0&-4\end{bmatrix}x_{k}\right\}u_{k}$ (12)
$=\begin{bmatrix}a_{1}&a_{2}\\ a_{4}&a_{5}\end{bmatrix}x_{k}+\left\{\begin{bmatrix}b_{1}\\ b_{2}\end{bmatrix}+\begin{bmatrix}a_{3}&0\\ 0&a_{6}\end{bmatrix}x_{k}\right\}u_{k},$ (13)